# Unraveling the Connections between Flow Matching and Diffusion Probabilistic Models in Training-free Conditional Generation

### v1

@@ Abstract
Training-free conditional generation aims to leverage the unconditional diffusion models to implement the conditional generation, where flow-matching (FM) and diffusion probabilistic models (DPMs) are two mature unconditional diffusion models that achieve high-quality generation. Two questions were asked in this paper: What are the underlying connections between FM and DPMs in training-free conditional generation? Can we leverage DPMs to improve the training-free conditional generation for FM? We first show that a probabilistic diffusion path can be associated with the FM and DPMs. Then, we reformulate the ordinary differential equation (ODE) of FM based on the score function of DPMs, and thus, the conditions in FM can be incorporated as those in DPMs. Finally, we propose two posterior sampling methods to estimate the conditional term and achieve a training-free conditional generation of FM. Experimental results show that our proposed method could be implemented for various conditional generation tasks. Our method can generate higher-quality results than the state-of-the-art methods.
Figure 1: Our proposed FMPS in pre-trained pixel and latent flow diffusion models in various conditional generation applications. FMPS has the same flexibility and general as the training-free methods for diffusion probabilistic models.
@@ 1 Introduction
Diffusion-based conditional generation methods [1, 2, 3] aims at generating desired images according to the given conditions, such as text prompts. Conditional generation is the key component for various computer vision tasks, such as linear/non-linear inverse problems [2, 1], image edit [4], etc. These methods can be roughly categorized into training-based [5, 6] and training-free [7, 3] approaches. In this paper, we mainly focus on the training-free method.
Training-free conditional generation methods [1, 3, 8] rely on a pre-trained unconditional diffusion model and do not require retraining. Based on the way to build the probabilistic path from Gaussian noise to data distribution, the unconditional diffusion models can be divided into two categories: diffusion probabilistic models (DPMs) [9, 10, 11], and flow diffusion models [12]. DPMs introduced the score function [9] to model the probabilistic path as a stochastic path. Flow diffusion models used Continuous Normalizing Flow (CNF) [13] to model the probabilistic path via flow-matching (FM), where CNF enables an interpolation way to define the probabilistic path from two distribution points. To simplify the notation, we abbreviated flow diffusion models as FMs to emphasize the flow-match.
When the unconditional diffusion model is a DPM, one of the training-free conditional DPMs is based on the posterior sampling [2]. DPS [2] first proposed the posterior sampling to solve the linear inverse problems. FreeDom [2] improved DPS by introducing the energy-based function to extend posterior sampling to solve non-linear inverse problems. MPGD [3] further introduced the manifold theory to improve FreeDom. The posterior sampling methods provide a flexible solution for various conditional generation tasks.
Unlike DPMs, there was very little research on training-free conditional FMs. The latest works include D-flow [14] and $\pi$-flow [15]. D-flow guided the generation process by fine-tuning the start point of FMs. $\pi$-flow [15] only focuses on the linear inverse problems for conditional generation.
The existing training-free methods for FMs have certain limitations, e.g., the $\pi$-flow is hidden on the linear inverse assumption. While the posterior sampling [2] in DPMs can incorporate the conditions into the unconditional diffusion model easily. Unlike DPMs, FM does not have a direct way of introducing the conditions. Therefore, two natural questions arise: 1) Is there any connections between FM and DPMs? and 2) Can we leverage posterior sampling in DPMs to improve the training-free conditional generation for FM?
In this paper, we propose a novel training-free conditional generation method for FMs called flow-match based posterior sampling (FMPS). The key insight is first to unravel the connection, where there is a probabilistic path between FMs and DPMs. This enables FMPS to define the score function explicitly. In this condition, FMPS could achieve the training-free conditional generation following a similar road of DPMs via posterior sampling.
Concretely, posterior sampling depends on the score function in DPMs. We first explicitly connect the ODE of FM and the score function in DPMs (Detailed in Sec. 4). Then, to make the posterior sampling available for FMs, we redefine the ODE of the FMs as the score function in DPMs. This new definition could introduce conditions into FMs via posterior sampling, where FMPS can also easily change the marginal distribution (i.e., the unconditional FMs) to the conditional distribution (i.e., the conditional FMs). Hence, an unconditional FM and a distance metric could be used for the training-free conditional FMs. Moreover, we also propose two posterior sampling methods: gradient-aware FMPS and gradient-free FMPS, where the difference is whether we need to calculate the gradient of the FMs to perform the posterior sampling. The experimental results on linear inverse problems, non-linear inverse problems, and text-to-image generation tasks show that FMPS achieves better generation quality in various downstream tasks and verify the validity and efficiency of FMPS.
@@
  • We unravel the connections between flow-matching and diffusion probabilistic models.
@@ 2 Related Work
Diffusion-based conditional generation tasks. Diffusion-based conditional generation tasks aim at leveraging the potential of the diffusion models [10, 11, 9] to generate the desired samples, which is an important challenge. Conditional generation has been widely applied in various ways. For example, the inverse problem [16, 2] uses degenerated images as a condition to enhance its quality. The text-to image [17, 18, 19, 20], text-to-video [21], and text-to-3D [22] generations use text prompts as a condition to generate the images, the videos, the 3D objects described by the text prompt, respectively. Depending on whether the condition generation model needs to be retrained, the conditional generation methods can be mainly divided into training-based and training-free methods. In this paper, we focus on the training-free methods.
Training-free methods [2, 1, 3, 16, 15] aim at leveraging pre-trained unconditional diffusion models. Since the unconditional diffusion models can be categorized into DPMs and FMs, the training-free conditional DPMs and the training-free conditional FMs are discussed.
Training-free conditional DPMs. These methods [2, 1, 3, 23] use DPMs as the unconditional diffusion models, where DPMs model the sampling process [11, 10, 9] based on the score function [9]. One of the successful approaches is based on the posterior sampling proposed by DPS [2]. DPS indicated that the score function represents the correlation between marginal distribution, i.e., the target of the unconditional generation and the sampling process. Thus, DPS derived a conditional term to connect the unconditional generation and condition, leveraging this correction. Meanwhile, DPS further proposed an estimation method for the conditional term under linear inverse assumption [16]. Then, FreeDom [1] introduced the energy function to estimate the conditional term without the limitation of linear inverse assumption. MPGD [3] and DSG [7] introduced the manifold theory and sphere constraint to enhance the FreeDom, respectively.
Training-free conditional FMs. Recently, FMs [12, 13] have been proposed to leverage an ODE to model the sampling process via flow-matching [12]. Many researches [18, 13] show the potential of FMs, which makes it a potential choice for training-free methods. Since there is no explicit score function, training-free methods for FMs aim at building the connection between the unconditional generation process and condition. $\pi$Flow first proposed a closed-form after introducing the condition into the flow path under the linear inverse assumption. D-Flow [14] fine-tuned the initialized Gaussian noise to build the connection condition and generated images.
However, training-free conditional FMs are not flexible compared to the DPMs. In this paper, we observe an implicit score function hidden in the ODE of FMs. Leveraging this, we explicitly define the score function-based ODE for FMs. After the score function is combined, we could achieve flexible generation via FMs, where the ability of the FMs bootstraps the generation quality.
@@ 3 Preliminary
3.1 Training-free conditional DPMs
Unconditional diffusion probabilistic models. It introduces the Gaussian probabilistic path to model the forward process of Eq. 7 as:
$x_{t}=\sqrt{\hat{\alpha}_{t}}x_{0}+(1-\hat{\alpha}_{t})\epsilon,$ (1)
@@ 3.1 Training-free conditional DPMs
Then, the reverse process of DPMs could be defined based on the score function:
@@
where $f(t)$ and $g(t)$ are also the parameter of noise scheduler related to the $\hat{\alpha}_{t}$. $\nabla_{x_{t}}\log p(x_{t})$ is the score function, where it can also be represented as the predicted noise formulation: $\epsilon_{\theta}(x_{t},t)$ and is modeled by the neural network.
Training-free conditional DPMs via unconditional DPMs. Given a condition $c$ such as a text prompt, we could change the unconditional score function to the condition distribution as:
@@
$\nabla_{x_{t}}\log p(x_{t})$ is exactly the unconditional score function, which can be used directly. But $\nabla_{x_{t}}\log p(c|x_{t})$ is the conditional term, which is unknown and needs to be estimated.
Posterior Sampling on DPMs. Previous works [1, 2] proposed posterior sampling for the estimation of $\nabla_{x_{t}}\log p(c|x_{t})$. In this paper, we follow Freedom [1], which indicates we could use a differentiate distance metric $D(\ast,\ast)$ to measure the distance between $x_{t}$ and $c$. That is, $\nabla_{x_{t}}\log p(c|x_{t})$ could be directly estimated by $D(\ast,\ast)$ as:
@@
where $r$ is the hyperparameter to control the strength of given condition $c$. $\hat{x}_{0|t}$ is the data distribution prediction for $x_{t}$, which is defined as:
$\hat{x}_{0|t}=\frac{1}{\hat{a}_{t}}(x_{t}-(1-\hat{a}_{t})\epsilon_{\theta}(x_{t},t)).$ (6)
Therefore, by Eq. 4 - Eq. 6, we could achieve the training-free conditional generation.
3.2 Flow diffusion models
Flow match-based diffusion models (FMs) define the forward process between data distribution $p_{0}$ and the Gaussian noise distribution $p_{1}$ as follows:
@@ 3.2 Flow diffusion models
Then, FMs define the reverse process based on the ODE as:
@@
where $v_{\theta}(x_{t},t)$ is called the velocity function, which is modeled by the neural networks.
Thus, the reverse process could be calculated by the ODE solver such as Euler solve:
$x_{t-1}=x_{t}+hv_{\theta}(x_{t},t),$ (9)
where $h$ is the time interval between $t$-th time step to $t-1$-th time step.
Motivation. Focusing on Eq. 7 and Eq. 1, we could find that there is an implicit score function hidden in the ODE of FMs. Since we could set $a_{t}=\sqrt{\hat{\alpha}_{t}}$ and $b_{t}=1-\hat{\alpha}_{t}$. Then, the path from the FMs and DPMs are the same. Thus, the ODE of FMs could be represented by the score function similar to the DPMs. If we can explicitly define the score function to re-formulate the ODE of FMs, posterior sampling could help FMs achieve conditional generation again!
@@ 4 Methodology
In this paper, we propose a novel training-free conditional generation method for FMs that leverages posterior sampling in DPMs, called flow-matching based posterior sampling (FMPS). The first step for FMPS is to re-formulate the ODE of the reverse process via the score function. This enables FMPS to leverage the Bayesian rule, the key step for posterior sampling, to introduce the condition $c$ by changing the marginal distribution to the conditional distribution. In the end, using our new conditional term-based reverse process, FMPS proposes two new methods to estimate it and achieve conditional generation.
Connections between FMs and DPMs in the training-free conditional generation. To re-formulate the ODE of the reverse process, we have the following proposition:

### v2

# Flow Matching Posterior Sampling: A Training-free Conditional Generation for Flow Matching

@@ Abstract
Training-free conditional generation based on flow matching aims to leverage pre-trained unconditional flow matching models to perform conditional generation without retraining. Recently, a successful training-free conditional generation approach incorporates conditions via posterior sampling, which relies on the availability of a score function in the unconditional diffusion model. However, flow matching models do not possess an explicit score function, rendering such a strategy inapplicable. Approximate posterior sampling for flow matching has been explored, but it is limited to linear inverse problems. In this paper, we propose Flow Matching-based Posterior Sampling (FMPS) to expand its application scope. We introduce a correction term by steering the velocity field. This correction term can be reformulated to incorporate a surrogate score function, thereby bridging the gap between flow matching models and score-based posterior sampling. Hence, FMPS enables the posterior sampling to be adjusted within the flow matching framework. Further, we propose two practical implementations of the correction mechanism: one aimed at improving generation quality, and the other focused on computational efficiency. Experimental results on diverse conditional generation tasks demonstrate that our method achieves superior generation quality compared to existing state-of-the-art approaches, validating the effectiveness and generality of FMPS.
Figure 1: Our proposed FMPS in various conditional generation applications. FMPS has the same flexibility and generality as the training-free methods based on SDMs.
@@ 1 Introduction
Flow matching models (FMs) [1] have achieved tremendous success in generative modeling. These models learn a velocity function [2] to move the Gaussian distribution to the target data distribution. Although FMs have achieved impressive results, these models are trained without conditions. When generating images or videos, users often prefer to create them according to specific requirements. Therefore, conditional generative diffusion models [3, 4] have also received increasing attention recently. This is an essential task in generative modeling, which has a wide range of applications, e.g., linear/non-linear inverse problems [5], image editing [6], etc.
A simple and effective approach to obtaining a conditional generative diffusion model is to leverage an already pre-trained unconditional diffusion model. Please note that Score function-based Diffusion Models (SDMs) [7] and FMs represent two major categories of the unconditional diffusion models. In this paper, we only focus on conditional generation based on flow matching.
To leverage an unconditional diffusion model for conditional generation, a straightforward approach is to fine-tune this unconditional model by incorporating the condition into its backbone architecture [8, 9]. For example, D-flow [10] guides the generation process by fine-tuning the start point of FMs. However, fine-tuning a diffusion model incurs significant computational costs [5].
To reduce this cost, training-free conditional generation methods [11] have been proposed, which simply reuse the unconditional FMs to perform the conditional generation without any fine-tuning or retraining. For example, in unconditional SDMs [7], one successful training-free conditional generation approach is the posterior sampling [5]. Posterior sampling leverages Bayes’ rule to transform the unconditioned marginal distribution $p(x)$ to the unknown but desired conditional distribution $p(x|c)$, utilizing the score function inherent to SDMs. Further, Freedom [12] proposed using a simple energy function in the conditional term by computing the gradient of a distance metric between the current state and the condition. In this way, the reverse process reduces the distance between the current state and the condition, thereby achieving conditional generation. Later, MPGD [13] has been proposed to reduce the computation cost of posterior sampling by leveraging manifold theory.
Posterior sampling enables a flexible mechanism for incorporating the conditions, as it only requires the definition of a differentiable distance metric between the current state and the condition. This flexibility allows training-free conditional generation based on SDMs to be applied to various conditional generative tasks, such as inversion tasks [14] and style transfer tasks [15].
However, FMs learn a velocity function and do not have an explicit score function. It hinders the training-free conditional FMs from using posterior sampling, as the bridge that connects the unconditional distribution $p(x)$ and the conditional distribution $p(x|c)$ is missing. Only a few works have attempted to rebuild this connection. For example, $\pi$-flow [16] proposed posterior sampling for FMs, but it is limited to linear inverse problems. Feng et al. [17] further introduced the mentor carol approximation to incorporate posterior sampling with the energy function, yet it also applies only to linear inverse problems. Thus, existing posterior sampling methods for training-free conditional FMs remain limited in scope and cannot support flexible designs for diverse conditional generation tasks.
To address this problem, we propose a novel training-free conditional generation method for FMs, termed flow-matching posterior sampling (FMPS). FMPS enables posterior sampling to be as flexible as in SDMs, thereby extending conditional FMs beyond the domain of linear inverse problems. Our key insight is to introduce a correction term by steering the velocity field, which constructs a valid pathway linking the velocity function to the score function. This enables FMPS to achieve the same flexibility as SDMs in handling conditional generation.
Specifically, to incorporate the condition into the velocity field of unconditional FMs, we propose a correction term that provides an additional direction aligned with the condition, i.e., steering the velocity field to the condition. We then reformulate this correction term to resemble the score function in SDMs. Under this formulation, the entire framework reduces to posterior sampling in SDMs, where the conditional distribution can be formulated as the unconditional distribution and the energy function.
Furthermore, we propose two variants of FMPS: gradient-aware FMPS and gradient-free FMPS. The gradient-aware FMPS improves generation quality by leveraging the gradient of the FMs, while the gradient-free FMPS is more computationally efficient by avoiding gradient computations. The experimental results on linear inverse problems, non-linear inverse problems, and text-to-image generation tasks demonstrate that FMPS achieves superior generation quality in various downstream tasks, verifying the validity and efficiency of FMPS.
@@
  • We propose two straightforward approaches for implementing posterior sampling for FMs.
@@ 2 Related Work
Diffusion-based conditional generation has become a key approach for generating desired samples using diffusion models [18, 7, 19], which has been widely explored in various applications. For instance, in inverse problems [20, 5, 21, 22], degraded images are used as conditions to restore or enhance image quality. In addition, text-to-image [3, 2, 23, 24], text-to-video [25], and text-to-3D [4] generation methods utilize text prompts as conditions to generate corresponding images, videos, and 3D objects, respectively. Depending on whether the model requires retraining for condition incorporation, existing conditional generation methods can be broadly categorized into training-based and training-free approaches. This paper focuses on the latter, training-free methods.
Training-free methods [5, 12, 13, 20, 16] aim to leverage pre-trained unconditional diffusion models without requiring additional training. Based on the type of pre-trained model, i.e., Score-based Diffusion Models (SDMs) [26] and Flow Matching models (FMs) [27, 28], training-free conditional generation methods can be categorized into two groups: those based on SDMs and those based on FMs. In the following, we first provide an overview of SDMs and FMs in Subsection 2.1. We then discuss training-free conditional generation methods based on SDMs in Subsection 2.2 and those based on FMs in Subsection 2.3.
2.1 Two Types of Diffusion Models
There are two broad categories of diffusion models: score function-based [7, 18] and flow-matching-based [1, 2] diffusion models. The key distinction lies in the learning objective. Score-based diffusion models (SDMs) aim to learn a score function, which represents the divergence at the current state. This divergence, defined as the gradient of the log marginal distribution, guides the diffusion process toward the target data distribution [7]. As a result, SDMs can be interpreted as sampling from a marginal distribution $p(x)$, a property that is crucial for training-free methods. In contrast, Flow matching models (FMs) learn a velocity field at each state [1], which characterizes the dynamics along a path connecting the target data and noise distributions. FMs utilize this velocity field to parameterize the transition path. Generally, a straighter path corresponds to faster convergence. Therefore, FMs can learn a near-straight path to enable both fast and high-quality sampling.
2.2 Training-free Conditional Generation based on SDMs
Most existing training-free conditional generation methods [5, 12, 13, 29, 30, 14] are based on the score function in SDMs. The score function enables models to sample from an unknown conditional distribution $p(x|c)$ via the known marginal distribution $p(x)$, i.e., the posterior sampling. The difference behind these methods is how to approximate $p(x|c)$ via $p(x)$. DPS [5] proposed the Bayesian approximation method for $p(x|c)$ by leveraging the linear assumption in linear inversion. FreeDom [12] extended DPS by introducing the energy function. Such a design unleashed the training-free methods since we could approximate the $p(x|c)$ by defining a differentiable distance metric. In this case, SDM-based training-free method could solve the non-linear inverse problems [12] and more complex tasks such as style-transfer [31]. DSG [32] proposed the sphere theory to improve the posterior sample, and LGD-MC [33] proposed a multi-sampling process to improve the posterior sample. However, posterior sampling required the gradient of the diffusion models, thus introducing a computational cost problem. MPGD [13] introduced the linear manifold theory to avoid gradient calculation, thereby alleviating the computation cost. UFG [34] propose a unified principle for SDMS-based methods to select the hyperparameter.
2.3 Training-free Conditional Generation based on FMs
The FM paradigm has been introduced to improve upon SDMs by replacing the score function with a velocity field. However, this change disrupts the previously established link to the conditional distribution $p(x|c)$, which was naturally accessible through the score function in SDMs. As a result, FMs face challenges in reconstructing an effective connection between the marginal distribution $p(x)$ and the conditional distribution $p(x|c)$.
Several approaches have been proposed to address this issue. For instance, D-Flow [10] introduced a lightweight framework that fine-tunes the initial point of the generation process using a given condition, effectively guiding the trajectory toward $p(x|c)$. In contrast, $\pi$Flow [16] first exploits the theoretical relationship between FMs and SDMs to re-establish a formulation of $p(x|c)$ based on $p(x)$, and then proposes a closed-form solution under the assumption of linear inversion. Feng et al. [17] developed a novel theoretical framework that directly links $p(x|c)$ with the distribution $p(x)$. This approach leads to a closed-form expression for $p(x|c)$, which, however, is analytically intractable. To overcome this limitation, the authors further apply the mentor-color theory to approximate the intractable form, again relying on the linear inversion assumption.
Unlike existing approaches that often depend on the linear inversion assumption, we propose FMPS, a flexible framework that can be applied to diverse conditional generation tasks without such restrictions.
@@ 3 Preliminary
3.1 Training-free conditional SDMs
Unconditional SDMs. In the forward process of diffusion models, SDMs define a Gaussian probabilistic path to add noise to the image as:
$x_{t}=\sqrt{\hat{\alpha}_{t}}x_{0}+\sqrt{(1-\hat{\alpha}_{t})}\epsilon,$ (1)
@@ 3.1 Training-free conditional SDMs
In the reverse process, SDMs learn to move the Gaussian data back to the target data via the score function as follows:
@@
where $f(t)$ and $g(t)$ are the hyper-parameters in the noise scheduler related to the $\hat{\alpha}_{t}$. $\nabla_{x_{t}}\log p(x_{t})$ is the score function modeled by the neural network. Eq. 2 is a divergence field that guides Gaussian noise flow to the data distribution.
Training-free conditional SDMs. Given a condition $c$ such as a text prompt, the conditional score function that represents the divergence of the condition distribution can be defined as:
@@
$\nabla_{x_{t}}\log p(x_{t})$ is exactly the score function. $\nabla_{x_{t}}\log p(c|x_{t})$ is the conditional term, which is unknown and needs to be estimated.
Posterior Sampling on SDMs. Previous works [12, 5] proposed posterior sampling for the estimation of $\nabla_{x_{t}}\log p(c|x_{t})$. In this paper, we follow Freedom [12], which demonstrates that we could use a differentiate distance metric $D(\ast,\ast)$ to measure the distance between $x_{t}$ and $c$. That is, $\nabla_{x_{t}}\log p(c|x_{t})$ could be directly estimated by $D(\ast,\ast)$ as:
@@
where $r$ is the hyperparameter to control the strength of the given condition $c$. $\hat{x}_{0|t}$ is the data distribution prediction for $x_{t}$, which is defined as:
$\hat{x}_{0|t}=\frac{1}{\sqrt{\hat{a}_{t}}}(x_{t}-(1-\hat{a}_{t})\epsilon_{\theta}(x_{t},t)).$ (6)
Therefore, by Eq. 4 - Eq. 6, we achieve the training-free conditional generation based on SDMs.
3.2 Flow Matching Models
In the forward process, FMs move the data distribution $p_{0}$ to the Gaussian noise distribution $p_{1}$ as follows:
@@ 3.2 Flow Matching Models
In the reverse process, the velocity function is first defined as:
@@
where $v_{\theta}(x_{t},t)$ is the velocity function, which is modeled by the neural networks. Then, the reverse process could be calculated by the ODE solver, such as the Euler solver:
$x_{t-1}=x_{t}+h\times v_{\theta}(x_{t},t),$ (9)
where $h$ is the time interval between the $t$-th and $(t-1)$-th time steps.
3.3 An Inherent Challenge for Conditional Guidance for FMs
There is no score function in FMs as shown in Eq. 8. In this condition, the posterior sampling is invalid because there is no direct way to formulate $\nabla_{x}\log p(c|x)$ term as in Eq. 4. Thus, we cannot ensure the reverse process is guided to the direction of the condition.
To tackle this, Feng et al. [17] proposed a framework to introduce the guidance term as follows:
$dx_{t}=v_{\theta}(x_{t},t)+g_{t}(x_{t})dt,$ (10)
where $v_{\theta}(x_{t},t)$ is the velocity function and the $g_{t}$ is the guidance term to model the direction for the condition. They showed that the energy function $J$ can be used to approximate $g_{t}(x_{t})$ via $p(x|c)\sim\frac{1}{Z}p(x)e^{-J(x,c)}$, where $Z$ is the normalization constant. The definition is:
[…]
where $z=(x_{0},\epsilon)$. Eq. 11 is intractable due to the need to calculate the guidance term to travel all possible data pairs $z$ [17]. Thus, $g_{t}(x_{t})$ is limited to the linear inverse problem.
In this paper, unlike Feng et al. [17] that used Monte Carlo sampling to approximate posterior sampling, we show that our method is a direct and straightforward closed-form approximation via the connection between SDMs and FMs.
@@ 4 Methodology
In this section, we introduce a simple and efficient method to calculate our correction term $\Delta(x_{t},c)$ via the posterior sampling. The key insight is to rebuild the connection between the reverse process of FMs and the score function. Leveraging this, FMPS contains two key steps: 1) We reformulate the velocity function in FMs as the score function in SDMs. This ensures the reverse process of FMS can be defined as the distribution $p(x)$ in SDMs. Thus, we can formulate our condition-related correction term $\Delta(x_{t},c)$. 2) We propose the FMPS that achieves the posterior sampling for FMs. This ensures FMPS can also sample the conditional distribution $p(x|c)$ via the correction term, thereby correctly achieving conditional generation. The existing posterior sampling theoretically proves the validity of FMPS.
4.1 The Condition-related Correction Term
Eq. 8 represents the velocity field connecting the data distribution and the Gaussian noise distribution, where $v_{\theta}(x_{t},t)$ is the direction to guide the movement of the $x_{t}$. To introduce the condition $c$ into the reverse process, we propose to add a new condition-related correction term into the velocity function as:
$dx_{t}=[v_{\theta}(x_{t},t)+r\Delta(x_{t},c)]dt,$ (12)
where $\Delta(x_{t},c)$ is the correction term to offer the additional direction to $c$. This way explicitly introduces the condition into the FMs. And $r$ is the scale to control the strength. The challenge is how to ensure $\Delta(x_{t},c)$ could offer a valid direction for the condition.
Score function ensures valid guidance direction. Similar to the score function, the valid guidance direction is that the correction term could also sample from $p(x)$ to $p(x|c)$ as in SDMs. To achieve it, we need to show that the velocity function in FMs can be reformulated as the score function in SDMs.
Firstly, we have the following proposition to introduce: